# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ II.3 Diagnosing Topological Phonons at Generic Momenta
In 2007, Fu and Kane (2007) proposed that the the calculation for topological invariant of $\mathbb{Z}_{2}$ topological insulators can be simplified in systems with both inversion ($\mathcal{P}$) and time-reversal ($\mathcal{T}$) symmetries. This method involves examining the parity of the electron occupation at time-reversal invariant momenta (TRIM). A significant contribution of this work was the illustration of how the utilization of symmetries can simplify the computation of topological invariant, thereby enhancing the understanding and analysis within the field. In conjunction with the “ten-fold way” theory discussed in Sec. II.1, which suggests that crystalline symmetries can give rise to novel topological phases, it becomes evident that symmetries play a pivotal role in simplifying the classification and diagnosis of topological states. This insight laid the groundwork for subsequent developments in symmetry-based indicators Po et al. (2017) and the independent advancement of topological quantum chemistry Bradlyn et al. (2017).
The symmetry-based indicator theory constitutes a conceptual framework that utilizes symmetries to simplify the calculation and classification of topological states across various symmetry classes, particularly within the context of the 230 space groups, and accounting for the presence or absence of time-reversal ($\mathcal{T}$) and SU(2) symmetries. The key idea of this theory is the construction of a gapped band structure space, denoted as {BS}, for a given specified band system within the realm of 230 space groups. This space {BS} is constructed from the symmetry attributes of the occupied bands at each high-symmetry momenta in the Brillouin zone:
@@
Figure 12 displays the nontrivial symmetry-based indicator groups $X_{\textrm{BS}}$ for the 230 space groups in the context of phonon spectra. The notation $\mathbb{Z}_{n}\times\mathbb{Z}_{m}$ signifies Abelian groups that are isomorphic to the respective quotient groups Po et al. (2017). Song et al. (2018); Zhang et al. (2020a); Zhang and Murakami (2021) linked the the symmetry-based indicator groups to the specific topological invariants based on rigorous derivations, providing both the general proof and the analytical expression to quantify these indicators. Specifically, a nontrivial symmetry-based indicator group (e.g., with nonzero values) may correspond to multiple distinct topological invariants, such as the Berry phase and the $\mathbb{Z}_{2}$-monopole charge, reflecting a one-to-many mapping inherent to indicator theories. In the subsequent sections, we will use space group #2, #3 and #82 as case studies to illustrate the implications of the symmetry-based indicators in diagnosing topological phonon.
Figure 13: In space group #2, which only preserves the inversion symmetry, the configuration of topological phonon bands is associated with the symmetry-based indicators. Yellow nodal rings with $z_{2,1}z_{2,2}z_{2,3}z_{4}=(0002)$ carry a nonzero $\mathbb{Z}_{2}$ monopole charge. Adapted from Ref. Song et al. (2018).
Figure 14: The symmetry-based indicator theory is employed to identify Weyl phonons in space groups #3 with $C_{2}$ symmetry and #82 with $S_{4}$ symmetry. The blue and yellow markers denote the positions of Weyl phonons with distinct chirality. This theoretical framework enables the diagnosis of Weyl phonons at the $k=0$ and $k=\pi$ planes in $C_{2}$-symmetric systems, and at the $k=0$ plane in $S_{4}$-symmetric systems. Adapted from Ref. Song et al. (2018).
II.3.1 Symmetry-based Indicators for Nodal-line Phonons
@@
where $N_{\pm}(\bm{K})$ represents the count of bands that exhibit even/odd parity under inversion symmetry at the four time-reversal-invariant momenta (TRIM) located on the $k_{i\in\{1,2,3\}}=\pi$ plane. The quantities $z_{2,1}z_{2,2}z_{2,3}z_{4}$ serve as the indicators for the symmetry-based indicator group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{4}$. The first three $\mathbb{Z}_{2}$ indicate the orientation and the location of the node-line/ring band crossings. Specifically, nodal lines are oriented along the reciprocal lattice direction $k_{i}$ when $z_{2,i}=1$, or they are situated around $k_{i}=\pi$. The final $z_{4}$ index indicates the count of nodal line/ring band crossings. An even number of nodal lines or rings within the Brillouin zone is signified by $z_{4}=0,2$, while an odd number is indicated by $z_{4}=1,3$. Combined the information from these four symmetry-based indicators, one can ascertain the entire configuration of nodal lines/rings throughout the BZ Zhang et al. (2019b, b, 2020a); Zhang and Murakami (2021); Zhang et al. (2022c).
@@
II.3.2 Symmetry-based Indicators for Weyl Phonons
@@
$\xi$ is the eigenvalue of the $C_{4}$ operator, the quantities $\omega_{2}^{0}$ and $\omega_{2}^{\pi}$ represent the counts of Weyl phonons located at the $k_{i}=0$ and $k_{i}=\pi$ planes, respectively. Since space group #82 constitutes a body-centered lattice, $\omega_{2}^{\pi}$ is null. Consequently, if $\omega_{2}^{0}$ equals 1, it signifies the presence of 4 modulo 8 Weyl phonons on the $k_{i}=0$ plane. This is illustrated in Fig. 14 (b), where the blue and yellow letters “w” denote the locations of Weyl phonons with distinct chirality Song et al. (2018); Zhang et al. (2020a); Zhang and Murakami (2021).
II.4 Comparison Between Different Topological Quasi-particles
Topological phonons, grounded in the principles of band theory, exhibit parallels with other topological quasiparticles, particularly those within spinless electronic band structures. They share characteristics such as various types of Weyl points, Dirac points, and nodal lines/rings, along with the emergence of topological surface states due to the “bulk-surface correspondence”. Nonetheless, the fundamental distinctions between phonons and electrons could pave the way for the revelation of novel topological phenomena unique to phonons. Moreover, topological phonons are endowed with distinctive attributes that set them apart from their counterparts among other topological quasiparticles.
For instance, in electronic systems, node-line or node-ring states can be destroyed by the spin-orbit coupling effect, which introduces a finite energy gap. In contrast, these states emerge in phonon spectra due to their inherent “zero spin” and “orbital-free” characteristics Zhang et al. (2019b). In three-dimensional systems, acoustic phonons inherently exhibit threefold degeneracy as Goldstone modes, and they transform into spin-1 Weyl phonons in chiral systems Zhang et al. (2018a). Phonons can also possess additional quantum degrees of freedom, such as pseudospin and angular momentum, which endow them with novel physical properties associated with their topological states Zhang et al. (2023b). The topological nature of phonons is inherently linked to the vibrational modes of the crystal lattice, leading to unique physical phenomena and practical applications Prodan and Prodan (2009b). The collective nature of lattice vibrations fosters interactions between phonons and other quasiparticles within the lattice, such as electrons or magnons, resulting in hybridized states and emergent phenomena that differentiate them from those observed solely in electronic or magnetic systems. In addition, topological phonons can be observed through various experimental techniques, including Raman scattering Zhang et al. (2023b), thermal transport Mangeolle et al. (2022), inelastic x-ray scattering Miao et al. (2018); Zhang et al. (2019b); Li et al. (2021a), neutron scattering Jin et al. (2022), and high-energy electron loss spectroscopy Li et al. (2023a), offering concrete evidence of their topological characteristics.
@@ III Examples of Topological Phonons: 1D and 2D Lattice Models
The band topology is closely related to the concept of geometric phase, first introduced by M. Berry in the context of quantum adiabatic evolution within arbitrary parameter space Berry (1984). In solid-state systems, the Bloch wave vector $\bm{k}$ offers a natural quantum adiabatic parameter space, and the geometric phase based on $\bm{k}$ was first introduced by J. Zak Zak (1989). Two different kinds of geometric phases can be defined according to different physical dimensions: (i) One is the line integral of Berry connection $\theta_{C}=\oint_{C}\mathrm{d}\bm{k}\cdot\bm{\mathbf{A}}(\bm{k})$, for any closed 1D loop $C$ within the Brillouin zone; (ii) The other is the surface integral of the Berry curvature $\theta_{S}=\iint_{S}\mathrm{d}\bm{S}\cdot\bm{\Omega}(\bm{k})$ for any 2D surface $S$. $\bm{\mathbf{A}}(\bm{k})=i\langle u_{\bm{k}}|\bm{\nabla_{k}}u_{\bm{k}}\rangle$ and $\bm{\Omega}(\bm{k})=\bm{\nabla_{k}}\times\bm{\mathbf{A}}$ are the Berry connection and curvature, respectively (with $|u_{\bm{k}}\rangle$ being the cell-periodic part of Bloch wave function). The former can be defined in 1D-3D but the latter exist only in 2D and 3D because the cross derivative is ill-defined in 1D.
In 1D systems, the integration of the Berry connection over the entire Brillouin zone yields a topological invariant known as the Zak phase Zak (1989); Xiao et al. (2010):
$\theta_{Z}=i\int^{\pi/a}_{-\pi/a}\mathrm{d}k\leavevmode\nobreak\ \left\langle u_{k}\Big{|}\frac{\partial u_{k}}{\partial k}\right\rangle,$ (53)
where $a$ is the lattice constant. The Zak phase, denoted as $\theta_{Z}$, is intricately linked to the electronic polarization, forming the cornerstone of the modern theory of electric charge polarization King-Smith and Vanderbilt (1993). Due to its oddness under inversion symmetry $\mathcal{P}$, the Zak phase assumes only two possible values, 0 or $\pi$, within systems possessing $\mathcal{P}$ symmetry. A $\theta_{Z}$ value of $\pi$ (or $\theta_{Z}=0$) signifies an inverted (or normal) band order, respectively (Fig. 15).
@@
In 2D systems, the integration of Berry curvature over the 2D BZ defines the topological Chern number Thouless et al. (1982a); Hasan and Kane (2010); Qi and Zhang (2011)
$C=\frac{1}{2\pi}\iint\mathrm{d}^{2}\bm{k}\leavevmode\nobreak\ \Omega_{z}(\bm{k})$ (54)
@@
In this section, we will introduce specific models that are used to realize topological phononic states in 1D and 2D systems. Possible materials realizations are discussed in the end.
@@ III.1 Topological Phonons in 1D: Su-Schrieffer-Heeger Model
Figure 17: In a diatomic chain with one vibrational degree of freedom per atom, the phononic Su-Schrieffer-Heeger (SSH) model exhibits distinct topological characteristics: (a) For a topologically trivial phonon band structure, the intra-cell force constant $K_{1}$ exceeds the inter-cell force constant $K_{2}$. This configuration yields a conventional band ordering where the out-of-phase mode (with parity $+$”) is higher in energy than the in-phase mode (with parity --”). (b) At the $X$ point, a degeneracy occurs between the in-phase and out-of-phase modes when $K_{1}$ is equal to $K_{2}$. (c) A topologically nontrivial phonon band emerges when $K_{1}$ is less than $K_{2}$, resulting in an inverted band order between the $\Gamma$ and $X$ points. (d) The eigenmode spectrum $\omega$ is plotted for a finite chain comprising 50 unit cells, varying as a function of $\Delta K/\bar{K}$ ($\Delta K=K_{2}-K_{1}$, $\bar{K}=(K_{2}+K_{1})/2$) under fixed boundary conditions. The reference frequency is given by $\omega_{0}=\sqrt{\frac{\bar{K}}{m}}$. The red and blue curves denote the topological end (edge) states, with their respective wave displacement fields illustrated in figures (e) and (f). Adapted from Ref. Liu et al. (2020b).
@@
$\begin{split}\theta_{Z}=&\int^{\frac{\pi}{a}}_{-\frac{\pi}{a}}\mathrm{d}k\leavevmode\nobreak\ i\langle u_{1}(k)|\partial_{k}u_{1}(k)\rangle\\ =&\left\{\begin{array}[]{cl}\pi&(K_{2}>K_{1})\\ 0&(K_{2}<K_{1})\end{array}\right.\end{split}$ (56)
The non-zero topological invariant $\theta_{Z}$ signifies the presence of edge states within the bulk band gap, as depicted in Fig. 17(d). The red and blue curves represent the in-gap edge states, which are localized at the left and right boundaries of the chain, respectively, as shown in Figs. 17(e)-(f). Notably, although such in-gap edge states often appear in various variants of the SSH model, they are not topological in general and their presence/absence depends on boundary conditions. Namely, this quantized Zak phase does have “bulk-boundary correspondence”. Only when the system has chiral symmetry, the bulk-boundary correspondence holds for the Zak phase, which is related to the winding number. The original SSH model for electrons has the chiral symmetry, but the phononic counterpart does not. However, in practice, in-gap edge states often emerge, in the cases of the nontrivial Zak phse, even when the chiral symmetry is absent. The phononic SSH model has been theoretically introduced in isostatic lattices Kane and Lubensky (2014) and has been experimentally observed in acoustic systems Xiao et al. (2015).
@@ III.2 Topological Phonons in 2D: Honeycomb Lattice Model
In his 1928 seminal work, P. Dirac introduced the relativistic equation for electrons, capable of describing both massive and massless particles. The equation is $i\partial_{t}|\psi(t)\rangle=H|\psi(t)\rangle$ with the Hamiltonian $H=m\beta+\sum^{d}_{i}p_{i}\alpha_{i}$ ($\hbar=c=1$) Dirac (1928). $m$ and $p_{i}$ are the mass and momentum, respectively, and $d$ refers to the spatial dimension. In order that the Dirac equation can reproduce the Klein-Gordon equation $(\bm{p}^{2}+m^{2}-\partial^{2}_{t})|\psi\rangle=0$, the introduced $\beta$ and $\alpha_{i}$ must satisfy $\beta^{2}=\alpha^{2}_{i}=1$, $\{\alpha_{i},\alpha_{j}\}=2\delta_{ij}$, and $\{\beta,\alpha_{i}\}=0$, which is the Clifford algebra. The anticommutation relations of $\alpha_{i}$ and $\beta$ indicate that $\alpha_{i}$ and $\beta$ must have matrix forms of which the smallest dimension is 2. In 2D case, i.e. $d=2$, the Dirac matrices $\alpha_{i}$ and $\beta$ have a simple representation: $\alpha_{x}=\sigma_{x}$, $\alpha_{y}=\sigma_{y}$, and $\beta=\sigma_{z}$ where $\sigma_{x,y,z}$ are Pauli matrices. Therefore, the 2D massive Hamiltonian is given as $H=k_{x}\sigma_{x}+k_{y}\sigma_{y}+m\sigma_{z}$. The case of $m=0$ describes the massless Dirac fermions in graphene Castro Neto et al. (2009) at the Brillouin corners $K$ and $K^{\prime}$. The cone-shaped band dispersion, stemming from the honeycomb lattice geometry, is not confined to spinless electronic states, prompting extensive exploration of topological physical phenomena in photonic crystals Wu and Hu (2015); Lu et al. (2014) and magnonic systems Zhang et al. (2013), and phonon systems Liu et al. (2020b).
@@
In electronic systems, the 2D Dirac semimetal with gapless Dirac cones is the parent phase for the QH family Liu et al. (2018, 2020b) which typically exist in honeycomb lattices. This principle can be extended to phononic systems as well. The 2D Dirac states can be constructed from stacked 1D SSH models, i.e. the brick lattice model, as shown in Fig. 18. The honeycomb lattice can be obtained by gradually stretching the brick lattice without closing or opening band gaps.
@@
Building upon the SSH model discussed earlier, the system at a constant $k_{y}$ exhibits a topologically nontrivial (trivial) band structure, characterized by an inverted (normal) band order along each $k_{x}$ line when $\tilde{K}>K$ ($\tilde{K}<K$), as shown in Fig. 17 (b). The case where $\tilde{K}=K$ represents a critical condition, resulting in a 1D band structure along $k_{x}$ with a vanishing gap at $k_{x}=\pi/a_{x}$, thereby forming a pair of Dirac cones at $(\pi/a_{x},\pm\pi/3a_{y})$, as shown in Fig. 17(c). These Dirac points are protected by $\mathcal{PT}$, ensuring that the Berry phase accumulated around any loop encircling a Dirac point is $\pi$, as evidenced by the abrupt discontinuity in the Zak phase $\theta^{x}_{Z}(k_{y})=\int^{2b_{x}}_{0}\mathrm{d}k_{x}\leavevmode\nobreak\ i\langle u_{1}(k_{x},k_{y})|\partial_{k_{x}}u_{1}(k_{x},k_{y})\rangle$ near $k_{y}=\pm\pi/3a_{y}$ [Fig. 17(d)]. The nonzero Zak phase can support localized edge modes, as depicted in the edge state spectrum of a 1D periodic ribbon of finite width along the $x$-direction in Fig. 17(e).
Figure 19: (a) Phononic model on a stretched brick lattice with lattice constants $\bm{a}_{1}=(a_{x},a_{y})=a(1+\cos\phi,\sin\phi)$, and $\bm{a}_{2}=(a_{x},-a_{y})$. Its reciprocal lattice vectors are $\bm{b}_{1}=(b_{x},b_{y})=(\pi/a_{x},\pi/a_{y})$ and $\bm{b}_{2}=(b_{x},-b_{y})$. The honeycomb lattice corresponds to the case of $\phi=\pi/3$. (b) 1D phononic band structure as function of $k_{x}$ for different $k_{y}$. The 1D phonon band is topologically nontrivial within the range $-b_{y}/3<k_{y}<b_{y}/3$, and it reverts to a trivial state outside this interval. (c) 2D phononic band showing a pair of Dirac cones at $(b_{x},\pm b_{y}/3)$. (d) Topological Zak phase $\theta^{x}_{Z}(k_{y})$ calculated as function of $k_{y}$ for each Wilson loop parallel to $k_{x}$ axis. (e) Phononic band structure of a system with finite width along the $y$-axis and periodicity along the $x$-axis. The red and blue curves denote the topological edge states.

### v2

@@ II.4 Diagnosing Topological Phonons at Generic Momenta
In 2007, fu2007topological proposed that the the calculation for topological invariant of $\mathbb{Z}_{2}$ topological insulators can be simplified in systems with both inversion ($\mathcal{P}$) and time-reversal ($\mathcal{T}$) symmetries. This method involves examining the parity of the electron occupation at time-reversal invariant momenta (TRIM). A significant contribution of this work was the illustration of how the utilization of symmetries can simplify the computation of topological invariant, thereby enhancing the understanding and analysis within the field. In conjunction with the “ten-fold way” theory discussed in Sec. II.2, which suggests that crystalline symmetries can give rise to novel topological phases, it becomes evident that symmetries play a pivotal role in simplifying the classification and diagnosis of topological states. This insight laid the groundwork for subsequent developments in symmetry-based indicators po2017symmetry and the independent advancement of topological quantum chemistry bradlyn2017topological .
The symmetry-based indicator theory constitutes a conceptual framework that utilizes symmetries to simplify the calculation and classification of topological states across various symmetry classes, particularly within the context of the 230 space groups, and accounting for the presence or absence of time-reversal ($\mathcal{T}$) and $\mathcal{T}^{2}=+1$. The key idea of this theory is the construction of a gapped band structure space, denoted as {BS}, for a given specified band system within the realm of 230 space groups. This space {BS} is constructed from the symmetry attributes of the occupied bands at each high-symmetry momenta in the Brillouin zone:
@@
Figure 4 displays the nontrivial symmetry-based indicator groups $X_{\textrm{BS}}$ for the 230 space groups in the context of phonon spectra. The notation $\mathbb{Z}_{n}\times\mathbb{Z}_{m}$ signifies Abelian groups that are isomorphic to the respective quotient groups po2017symmetry . song2018diagnosis ; zhang2020diagnosis ; zhang2021predicting linked the the symmetry-based indicator groups to the specific topological invariants based on rigorous derivations, providing both the general proof and the analytical expression to quantify these indicators. Specifically, a nontrivial symmetry-based indicator group (e.g., with nonzero values) may correspond to multiple distinct topological invariants, such as the Berry phase and the $\mathbb{Z}_{2}$-monopole charge, reflecting a one-to-many mapping inherent to indicator theories. In the subsequent sections, we will use space group #2, #3 and #82 as case studies to illustrate the implications of the symmetry-based indicators in diagnosing topological phonon.
Figure 10: In space group #2, which only preserves the inversion symmetry, the configuration of topological phonon bands is associated with the symmetry-based indicators. Yellow nodal rings with $z_{2,1}z_{2,2}z_{2,3}z_{4}=(0002)$ carry a nonzero $\mathbb{Z}_{2}$ monopole charge. Adapted from Ref. song2018diagnosis .
Figure 11: The symmetry-based indicator theory is employed to identify Weyl phonons in space groups #3 with $C_{2}$ symmetry and #82 with $S_{4}$ symmetry. The blue and yellow markers denote the positions of Weyl phonons with distinct chirality. This theoretical framework enables the diagnosis of Weyl phonons at the $k=0$ and $k=\pi$ planes in $C_{2}$-symmetric systems, and at the $k=0$ plane in $S_{4}$-symmetric systems. Adapted from Ref. song2018diagnosis .
II.4.1 Symmetry-based Indicators for Nodal-line Phonons
@@
where $N_{\pm}(\bm{K})$ represents the count of bands that exhibit even/odd parity under inversion symmetry at the four time-reversal-invariant momenta (TRIM) located on the $k_{i\in\{1,2,3\}}=\pi$ plane. The quantities $z_{2,1}z_{2,2}z_{2,3}z_{4}$ serve as the indicators for the symmetry-based indicator group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{4}$. The first three $\mathbb{Z}_{2}$ indicate the orientation and the location of the node-line/ring band crossings. Specifically, nodal lines are oriented along the reciprocal lattice direction $k_{i}$ when $z_{2,i}=1$, or they are situated around $k_{i}=\pi$. The final $z_{4}$ index indicates the count of nodal line/ring band crossings. An even number of nodal lines or rings within the Brillouin zone is signified by $z_{4}=0,2$, while an odd number is indicated by $z_{4}=1,3$. Combined the information from these four symmetry-based indicators, one can ascertain the entire configuration of nodal lines/rings throughout the BZ zhang2019phononic ; zhang2019phononic ; zhang2020diagnosis ; zhang2021predicting ; zhang2022endless .
@@
II.4.2 Symmetry-based Indicators for Weyl Phonons
@@
$\xi$ is the eigenvalue of the $C_{4}$ operator, the quantities $\omega_{2}^{0}$ and $\omega_{2}^{\pi}$ represent the counts of Weyl phonons located at the $k_{i}=0$ and $k_{i}=\pi$ planes, respectively. Since space group #82 constitutes a body-centered lattice, $\omega_{2}^{\pi}$ is null. Consequently, if $\omega_{2}^{0}$ equals 1, it signifies the presence of 4 modulo 8 Weyl phonons on the $k_{i}=0$ plane. This is illustrated in Fig. 11 (b), where the blue and yellow letters “w” denote the locations of Weyl phonons with distinct chirality song2018diagnosis ; zhang2020diagnosis ; zhang2021predicting .
@@ III Examples of Topological Phonons: 1D and 2D Lattice Models
The band topology is closely related to the concept of geometric phase, first introduced by M. Berry in the context of quantum adiabatic evolution within arbitrary parameter space Berry1984Mar . In solid-state systems, the Bloch wave vector $\bm{k}$ offers a natural quantum adiabatic parameter space, and the geometric phase based on $\bm{k}$ was first introduced by J. Zak Zak1989Jun . Two different kinds of geometric phases can be defined according to different physical dimensions: (i) One is the line integral of Berry connection $\theta_{C}=\oint_{C}\mathrm{d}\bm{k}\cdot\bm{\mathbf{A}}(\bm{k})$, for any closed 1D loop $C$ within the Brillouin zone; (ii) The other is the surface integral of the Berry curvature $\theta_{S}=\iint_{S}\mathrm{d}\bm{S}\cdot\bm{\Omega}(\bm{k})$ for any 2D surface $S$. $\bm{\mathbf{A}}(\bm{k})=i\langle u_{\bm{k}}|\bm{\nabla_{k}}u_{\bm{k}}\rangle$ and $\bm{\Omega}(\bm{k})=\bm{\nabla_{k}}\times\bm{\mathbf{A}}$ are the Berry connection and curvature, respectively (with $|u_{\bm{k}}\rangle$ being the cell-periodic part of Bloch wave function). The former can be defined in 1D-3D but the latter exist only in 2D and 3D because the cross derivative is ill-defined in 1D.
In 1D systems, the integration of the Berry connection over the entire Brillouin zone yields a topological invariant known as the Zak phase Zak1989Jun ; Xiao2010Jul :
$\theta_{Z}=i\int^{\pi/a}_{-\pi/a}\mathrm{d}k\penalty 10000\ \left\langle u_{k}\Big|\frac{\partial u_{k}}{\partial k}\right\rangle,$ (58)
where $a$ is the lattice constant. The Zak phase, denoted as $\theta_{Z}$, is intricately linked to the electronic polarization, forming the cornerstone of the modern theory of electric charge polarization King-Smith1993Jan . Due to its oddness under inversion symmetry $\mathcal{P}$, the Zak phase assumes only two possible values, 0 or $\pi$, within systems possessing $\mathcal{P}$ symmetry. A $\theta_{Z}$ value of $\pi$ (or $\theta_{Z}=0$) signifies an inverted (or normal) band order, respectively (Fig. 12).
@@
In 2D systems, the integration of Berry curvature over the 2D BZ defines the topological Chern number Thouless1982Aug ; hasan2010colloquium ; qi2011topological
$C=\frac{1}{2\pi}\iint\mathrm{d}^{2}\bm{k}\penalty 10000\ \Omega_{z}(\bm{k})$ (59)
@@
In this section, we will introduce specific models that are used to realize topological phononic states in 1D and 2D systems. Possible materials realizations are discussed in the end. The theoretical 1D and 2D models described in this section equally apply to crystalline solids and some metamaterials which have been theoretically proposed kane2014topological ; Jin2018Dec ; Li2020Feb and experimentally confirmed lu2018valley ; xiao2015geometric ; li2023direct .
@@ III.1 Topological Phonons in 1D: Su-Schrieffer-Heeger Model
Figure 14: In a diatomic chain with one vibrational degree of freedom per atom, the phononic Su-Schrieffer-Heeger (SSH) model exhibits distinct topological characteristics: (a) For a topologically trivial phonon band structure, the intra-cell force constant $K_{1}$ exceeds the inter-cell force constant $K_{2}$. This configuration yields a conventional band ordering where the out-of-phase mode (with parity ++”) is higher in energy than the in-phase mode (with parity -”). (b) At the $X$ point, a degeneracy occurs between the in-phase and out-of-phase modes when $K_{1}$ is equal to $K_{2}$. (c) A topologically nontrivial phonon band emerges when $K_{1}$ is less than $K_{2}$, resulting in an inverted band order between the $\Gamma$ and $X$ points. (d) The eigenmode spectrum $\omega$ is plotted for a finite chain comprising 50 unit cells, varying as a function of $\Delta K/\bar{K}$ ($\Delta K=K_{2}-K_{1}$, $\bar{K}=(K_{2}+K_{1})/2$) under fixed boundary conditions. The reference frequency is given by $\omega_{0}=\sqrt{\frac{\bar{K}}{m}}$. The red and blue curves denote the topological end (edge) states, with their respective wave displacement fields illustrated in figures (e) and (f). Adapted from Ref. liu2020topological .
@@
[…]
The non-zero topological invariant $\theta_{Z}$ signifies the presence of edge states within the bulk band gap, as depicted in Fig. 14(d). The red and blue curves represent the in-gap edge states, which are localized at the left and right boundaries of the chain, respectively, as shown in Figs. 14(e)-(f). Notably, although such in-gap edge states often appear in various variants of the SSH model, they are not topological in general and their presence/absence depends on boundary conditions. Namely, this quantized Zak phase does have “bulk-boundary correspondence”. Only when the system has chiral symmetry, the bulk-boundary correspondence holds for the Zak phase, which is related to the winding number. The original SSH model for electrons has the chiral symmetry, but the phononic counterpart does not. However, in practice, in-gap edge states often emerge, in the cases of the nontrivial Zak phse, even when the chiral symmetry is absent. The phononic SSH model has been theoretically introduced in isostatic lattices kane2014topological and has been experimentally observed in acoustic systems xiao2015geometric .
@@ III.2 Topological Phonons in 2D: Honeycomb Lattice Model
In his 1928 seminal work, P. Dirac introduced the relativistic equation for electrons, capable of describing both massive and massless particles. The equation is $i\partial_{t}|\psi(t)\rangle=H|\psi(t)\rangle$ with the Hamiltonian $H=m\beta+\sum^{d}_{i}p_{i}\alpha_{i}$ ($\hbar=c=1$) Dirac1928Feb . $m$ and $p_{i}$ are the mass and momentum, respectively, and $d$ refers to the spatial dimension. In order that the Dirac equation can reproduce the Klein-Gordon equation $(\bm{p}^{2}+m^{2}-\partial^{2}_{t})|\psi\rangle=0$, the introduced $\beta$ and $\alpha_{i}$ must satisfy $\beta^{2}=\alpha^{2}_{i}=1$, $\{\alpha_{i},\alpha_{j}\}=2\delta_{ij}$, and $\{\beta,\alpha_{i}\}=0$, which is the Clifford algebra. The anticommutation relations of $\alpha_{i}$ and $\beta$ indicate that $\alpha_{i}$ and $\beta$ must have matrix forms of which the smallest dimension is 2. In 2D case, i.e. $d=2$, the Dirac matrices $\alpha_{i}$ and $\beta$ have a simple representation: $\alpha_{x}=\sigma_{x}$, $\alpha_{y}=\sigma_{y}$, and $\beta=\sigma_{z}$ where $\sigma_{x,y,z}$ are Pauli matrices. Therefore, the 2D massive Hamiltonian is given as $H=k_{x}\sigma_{x}+k_{y}\sigma_{y}+m\sigma_{z}$. The case of $m=0$ describes the massless Dirac fermions in graphene castro2009electronic at the Brillouin corners $K$ and $K^{\prime}$. The cone-shaped band dispersion, stemming from the honeycomb lattice geometry, is not confined to spinless electronic states, prompting extensive exploration of topological physical phenomena in photonic crystals wu2015scheme ; lu2014topological and magnonic systems zhang2013topological , and phonon systems liu2020topological .
@@
In electronic systems, the 2D Dirac semimetal with gapless Dirac cones is the parent phase for the QH family liu2018berry ; liu2020topological which typically exist in honeycomb lattices. This principle can be extended to phononic systems as well. The 2D Dirac states can be constructed from stacked 1D SSH models, i.e. the brick lattice model, as shown in Fig. 15. The honeycomb lattice can be obtained by gradually stretching the brick lattice without closing or opening band gaps.
@@
Building upon the SSH model discussed earlier, the system at a constant $k_{y}$ exhibits a topologically nontrivial (trivial) band structure, characterized by an inverted (normal) band order along each $k_{x}$ line when $\tilde{K}>K$ ($\tilde{K}<K$), as shown in Fig. 14 (b). The case where $\tilde{K}=K$ represents a critical condition, resulting in a 1D band structure along $k_{x}$ with a vanishing gap at $k_{x}=\pi/a_{x}$, thereby forming a pair of Dirac cones at $(\pi/a_{x},\pm\pi/3a_{y})$, as shown in Fig. 14(c). These Dirac points are protected by $\mathcal{PT}$, ensuring that the Berry phase accumulated around any loop encircling a Dirac point is $\pi$, as evidenced by the abrupt discontinuity in the Zak phase $\theta^{x}_{Z}(k_{y})=\int^{2b_{x}}_{0}\mathrm{d}k_{x}\penalty 10000\ i\langle u_{1}(k_{x},k_{y})|\partial_{k_{x}}u_{1}(k_{x},k_{y})\rangle$ near $k_{y}=\pm\pi/3a_{y}$ [Fig. 14(d)]. The nonzero Zak phase can support localized edge modes, as depicted in the edge state spectrum of a 1D periodic ribbon of finite width along the $x$-direction in Fig. 14(e).
Figure 16: (a) Phononic model on a stretched brick lattice with lattice constants $\bm{a}_{1}=(a_{x},a_{y})=a(1+\cos\phi,\sin\phi)$, and $\bm{a}_{2}=(a_{x},-a_{y})$. Its reciprocal lattice vectors are $\bm{b}_{1}=(b_{x},b_{y})=(\pi/a_{x},\pi/a_{y})$ and $\bm{b}_{2}=(b_{x},-b_{y})$. The honeycomb lattice corresponds to the case of $\phi=\pi/3$. (b) 1D phononic band structure as function of $k_{x}$ for different $k_{y}$. The 1D phonon band is topologically nontrivial within the range $b_{y}/3<k_{y}<b_{y}/3$, and it reverts to a trivial state outside this interval. (c) 2D phononic band showing a pair of Dirac cones at $(b_{x},\pm b_{y}/3)$. (d) Topological Zak phase $\theta^{x}_{Z}(k_{y})$ calculated as function of $k_{y}$ for each Wilson loop parallel to $k_{x}$ axis. (e) Phononic band structure of a system with finite width along the $y$-axis and periodicity along the $x$-axis. The red and blue curves denote the topological edge states.